OpenStudy (anonymous):

A glass vase is shaped like a cylinder and has a radius of 2.5 inches. The vase holds 170 cubic inches of water. How tall is the vase? Use 3.14 for pi. Round your answer to the nearest hundredth. 8.66 inches 9.78 inches 10.35 inches 12.50 inches

OpenStudy (anonymous):

@satellite73 @mony01

OpenStudy (anonymous):

@roz_mar23@iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

h = V/(3.14r²)

OpenStudy (anonymous):

H=V/(3.14R^2)

OpenStudy (anonymous):

H=170/(3.14*2.5^2)

OpenStudy (anonymous):

@Synquence did u get it?

OpenStudy (anonymous):

No @roz_mar23

OpenStudy (anonymous):

The formula for height is H=V/(3.14R^2) Plug in the radius for R and volume of water for V

OpenStudy (anonymous):

Oh okay, one sec

OpenStudy (igreen):

Volume of a Cylinder: \(V = \pi r^2 h\)

OpenStudy (anonymous):

Thanks also mate @iGreen i got it 8.66

OpenStudy (igreen):

We are given: Radius = 2.5 Volume = 170 Now let's solve for \(h\). \(V = \pi r^2 h\) Plug in what we know: \(170 = (3.14) (2.5)^2 h\) Simplify Exponent: \(170 = (3.14) (6.25) h\) Multiply: \(170 = 19.625 h\) Divide 19.625 to both sides, what's 170 / 19.625? @Synquence

OpenStudy (igreen):

Yep, that's right. @Synquence